Ella looks at her digital clock. She sees that the number of minutes is 2 times the number of hours. The sum of the digit is 15. How do you solve it?
Answer by josgarithmetic(39617)

Assuming 12 hour format, the values would form H:M for hours on the left of the colon and minutes on the right side. The maximum value on the left would be 12. She sees M=2H.

If she saw the maximum H=12, then the corresponding M=2*12=24.
H_____________M=2*H
12____________24
11____________22
10____________20
09____________18
08____________16-----------That gives sum of the digits to 15.
----------------------------------8+1+6=15


That is crude, but it works.
The digital clock shows 08:16.